(a) Find the elasticity, using calculus, of P = -3Q + 18 when Q = 4

(b) Find the elasticity, using calculus, of P = Q^2-8Q+16 when Q=2
Answer #1

Here P = -30 to 0 30 ²18-8 Q = 6-P When a=4 P2 - 3x4 +18 in 6 PED = d (d) + b de do 6-f) dpt 4 PED = -0.5] Cinelastic demand)